OpenStudy (anonymous):

How do you solve -12x=72

OpenStudy (aroub):

x=-6

OpenStudy (aroub):

how.. u just divide by 12 both sides to get rid of it like this ----> -12*1/12 x=-72*1/12 x= -6

OpenStudy (aroub):

thank you!! :)

OpenStudy (angela210793):

always welcome :D

OpenStudy (anonymous):

divide both sides byt -12 so u get x = -6